# Air Quality Prediction with Physics-Informed Dual Neural ODEs in Open Systems

### v1

@@ Abstract
Air pollution significantly threatens human health and ecosystems, necessitating effective air quality prediction to inform public policy. Traditional approaches are generally categorized into physics-based and data-driven models. Physics-based models usually struggle with high computational demands and closed-system assumptions, while data-driven models may overlook essential physical dynamics, confusing the capturing of spatiotemporal correlations. Although some physics-informed approaches combine the strengths of both models, they often face a mismatch between explicit physical equations and implicit learned representations. To address these challenges, we propose Air-DualODE, a novel physics-informed approach that integrates dual branches of Neural ODEs for air quality prediction. The first branch applies open-system physical equations to capture spatiotemporal dependencies for learning physics dynamics, while the second branch identifies the dependencies not addressed by the first in a fully data-driven way. These dual representations are temporally aligned and fused to enhance prediction accuracy. Our experimental results demonstrate that Air-DualODE achieves state-of-the-art performance in predicting pollutant concentrations across various spatial scales, thereby offering a promising solution for real-world air quality challenges.
@@ 1 Introduction
Traditional research on air quality prediction largely relies on physics-based and data-driven approaches, each of which has their own inherent limitations. Physics-based models make predictions by solving Ordinary or Partial Differential Equations (ODEs/PDEs), which capture the transport of pollutants through diffusion and advection, as illustrated in Fig.1. Although these models achieve high accuracy, they require substantial computational resources, especially when applied on large spatial scales. Conversely, data-driven models excel at uncovering dependencies within data (Zhao et al., 2023; Cheng et al., 2023), yet their lack of integration with physical principles can lead to incomplete or even erroneous representations of spatiotemporal relationships. Therefore, it is crucial to propose a hybrid model that effectively harnesses the strengths of both physics-based and data-driven approaches. However, the development of such a model presents two primary challenges.
@@
In this paper, we propose Air-DualODE, a Physics-informed Dual Neural ODEs that integrate physics dynamics and data-driven dynamics for open system’s Air Pollution Prediction. For the first challenge, we propose a discrete Boundary-Aware Diffusion-Advection Equation (BA-DAE) to appropriately model open air systems. Due to the uncertainties associated with pollutant generation (source) and dissipation (sink) in open air systems, pollutant changes involve significant uncertainty, but both processes are linked to individual stations. To capture these dynamics, BA-DAE introduces a linear correction term, breaking away from the unrealistic closed system assumption and providing a more accurate description of physical phenomena in open air systems. Additionally, we incorporate Data-Driven Dynamics to capture spatiotemporal dependencies that BA-DAE ignores, making our approach better suited for real-world open air systems. For example, Data-Driven Dynamics can learn some patterns of pollutants generation, which cannot be described by BA-DAE.
@@ 2.3 Related Work
Air Quality Prediction. Air quality prediction is a key objective in smart city. Existing approaches can be broadly categorized into two types: physics-based models and deep learning models. The former rely on domain knowledge to construct Ordinary and Partial Differential Equations (ODEs/PDEs) that describe air pollutant transport (Li et al., 2023; Daly & Zannetti, 2007). However, these models face two major limitations. They rely on closed-system assumptions, and their use of numerical methods, such as finite difference or finite element methods (Özişik et al., 2017; Ŝolín, 2005), becomes computationally expensive at larger spatial scales. This makes rapid prediction difficult for urban planning and policy-making. Deep learning methods (Campos et al., 2023; 2024; Qiu et al., 2024; Chen et al., 2024) have recently gained prominence by using historical data and auxiliary covariates, such as weather, to extract high-dimensional spatiotemporal representations for pollutant prediction. These models typically employ RNNs or Transformers for temporal representations and GNNs or CNNs for spatial dependencies (Yu et al., 2017; Shang et al., 2021; Zhao et al., 2023; Wu et al., 2024). Furthermore, there are models specifically designed for air pollutant prediction tasks (Yu et al., 2023; Liang et al., 2023). However, these deep learning approaches may overlook essential physical dynamics, leading to confused spatiotemporal correlations that could violate physical principles.
Physics Guided Deep Learning. Recently, several studies have explored incorporating physical knowledge into deep learning models (Karpatne et al., 2024). The most direct approach involves using physical knowledge as constraints in loss functions (Raissi et al., 2017; Shi et al., 2021), penalizing deviations from physical laws to encourage physically consistent patterns and improve generalizability (Greydanus et al., 2019). However, these methods rely on highly accurate physical knowledge; otherwise, they may introduce inductive bias, limiting the model’s learning capacity. Another approach is to design hybrid frameworks that integrate scientific knowledge into specific modeling components (Bao et al., 2021; Xu et al., 2024), such as incorporating physical equations into Neural ODEs (Ji et al., 2022; Verma et al., 2024; Hettige et al., 2024). These methods often require deep latent representations to align with physical meanings, forcing models to switch between abstract representations and physical quantities Nascimento et al. (2021).
@@ 3.2.2 Discrete Boundary-Aware Diffusion-Advection Equation for open system
In an open system, changes in the total amount of pollutants can be categorized into dissipation and generation. Pollutant dissipation refers to sinks, including transport out of the boundary through diffusion and advection, as well as absorption by forests or lakes. Pollutant generation refers to sources such as industrial emissions. As shown in Fig.4, the dissipation and generation at Point E are highly relevant to its current concentrations. Therefore, we incorporate two terms into Eq.3: the dissipation term $\beta_{1}X$ ($\beta_{1}\in[-1,0]$) and the generation term $\beta_{2}X$ ($\beta_{2}\in[0,+\infty)$). These terms can be simplified into an open system correction term: $\beta X$ ($\beta\in[-1,+\infty)$). Boundary stations, experiencing more outward diffusion and advection, tend to have greater negative $\beta$ values. So, this equation is called as Discrete Boundary-Aware Diffusion-Advection Equation (BA-DAE):
$\mathbf{F^{P}}(X_{t};\mathbf{\Theta})=\frac{dX}{dt}=\alpha\odot(-k\cdot L_{\text{diff}}X)+(1-\alpha)\odot(L_{\text{adv}}X)+\beta X.$ (6)
Here, $\mathbf{\Theta}$ represents learnable parameters used to approximate the graph Laplacian operators. The gate value $\alpha\in\mathbb{R}^{N\times 1}$ is estimated by a linear layer ($\mathbb{R}^{2}\to\mathbb{R}$) based on each station’s diffusion and advection over time, representing the weighted influence of these two effects as they vary across time and space. The diffusion coefficient $k$ and the correction term’s coefficient $\beta$ are estimated via the Coefficient Estimator, a Recurrent Neural Network. By solving BA-DAE by applying ODE solver from Chen et al. (2018), pollutant concentrations from $T+1$ to $T+\tau$ are obtained, and a temporal encoder maps these concentrations to the latent space for dynamics fusion, as illustrated in the Physics Dynamics section of Fig 2.
@@ 3.3 Data-Driven Neural ODE
where $\mathbf{\Phi}$ represents learnable parameters in ODE function $\mathbf{F^{D}}$. To capture spatial correlations across different time steps within the dynamic system, a masked self-attention mechanism is incorporated into $\mathbf{F^{D}}$. Specifically, the adjacency matrix of $G$ is employed as a mask in the Spatial-MSA since intuitively, if no potential transport pathway exists between two stations, their representations should not be correlated. This approach not only improves computational efficiency but also enables the model to focus on relevant information, enhancing its effectiveness across various spatial granularities in prediction scenarios.
@@ 3.4 Dynamics Fusion
Decaying Temporal Constrastive Learning. Before combining the physics and data-driven dynamics, it is essential to align their latent representations, $Z^{P}_{T+1:T+\tau}$ and $Z^{D}_{T+1:T+\tau}$, to resolve inconsistencies and enable effective fusion. Since both dynamics operate synchronously, three key intuitions are considered: 1) latent representations at the same time step are more similar than those at different time steps, 2) adjacent time steps are more similar than distant ones, and 3) representations from the same dynamics are more similar than those from different dynamics. To satisfy the three intuitions, we propose Decaying Temporal Contrastive Learning (Decay-TCL). It computes a decaying weight $w(t,s)$ based on timestamp differences and dynamics type to regulate the similarity between latent representations. The hyperparameters $\lambda_{1}$ and $\lambda_{2}$ are used to differentiate the distinct dynamics.
@@
GNN Fusion. After temporal alignment, the two latent representations are concatenated as $Z^{(0)}_{t}=\text{concat}(Z^{P}_{t},Z^{D}_{t})$ at each timestamp. Since the influence of stations is distance-dependent, using Attention and MLP structures may introduce noise from distant sites, complicating training. To mitigate this, a GNN is employed based on the geospatial graph $G$ to enhance Dynamics Fusion: $Z^{(n)}_{t}=\text{GNN}(Z^{(0)}_{t})$, where $n$ represents the GNN layers. The Decoder then transforms the latent representations $Z^{(n)}_{T+1:T+\tau}$ into predicted pollutant concentrations $\hat{X}_{T+1:T+\tau}$. The prediction loss is computed as follows:
$\mathcal{L}_{\text{pred}}=\frac{1}{\tau}\sum_{t=T+1}^{T+\tau}\|X_{t}-\hat{X}_{t}\|.$ (13)
We jointly optimize the loss functions in Eq.12 and Eq.13, with a hyperparameter $\gamma$ to balance them.
$\mathcal{L}=\mathcal{L}_{\text{pred}}+\gamma\mathcal{L}_{\text{tcl}}.$ (14)
@@ 4.1 Experiment Setting
Baselines. We evaluate Air-DualODE against baselines of four categories: 1) Classical Methods: Historical Average (HA) and Vector Auto-Regression (VAR). 2) Differential Equation Network-Based Methods: Latent-ODE  (Chen et al., 2018), ODE-RNN (Rubanova et al., 2019) and ODE-LSTM  (Lechner & Hasani, 2020). 3) Spatio-temporal Deep Learning Methods: Diffusion Convolution Recurrent Neural Network (DCRNN)  (Li et al., 2017), Spatio-temporal Graph Convolutional Network (STGCN)  (Yu et al., 2017), ASTGCN  (Guo et al., 2019), GTS  (Shang et al., 2021), PM2.5-GNN  (Wang et al., 2020), and AirFormer  (Liang et al., 2023). 4) Physics-informed Neural Networks: AirPhyNet (Hettige et al., 2024).
@@ 4.2 Performance Comparison
From Table 1 we can also observe the following: 1) deep learning approaches outperform classical methods such as HA and VAR, demonstrating their superior ability to capture complex spatiotemporal correlations. 2) Spatiotemporal deep learning models, originally developed for traffic flow forecasting, such as ASTGCN and GTS, exhibit strong adaptability to air quality forecasting, producing results comparable to SOTA air quality prediction methods like AirFormer and PM25GNN. 3) Approaches based on Neural DE networks also exhibit competitive performance compared to spatiotemporal deep learning methods, as they capture the ODE function in latent space, which is crucial for accurate air quality prediction. Overall, these results suggest the effectiveness of Air-DualODE in achieving precise air quality predictions while leveraging the physics-based domain knowledge of air pollutant transport.
@@ 4.3 Ablation Study
Effect of Physical knowledge. To further demonstrate the impact of incorporating physical knowledge, we compare Air-DualODE with three other closed-system variations: a) Diff-c, which only considers the diffusion process, b) Adv-c, which only considers the advection process, and c) DAE-c, which accounts for both processes. Additionally, we compare these with d) BA-DAE-o, namely Air-DualODE, as proposed in Section 3.2.2 for open air systems. Fig.5 shows that integrated physical knowledge is crucial, and inadequate modeling leads to reduced performance. Consistent with AirPhyNet (Hettige et al., 2024), DAE-c outperforms Diff-c and Adv-c, but its inability to model dissipation and generation makes it less effective than BA-DAE-o. This underscores the effectiveness of BA-DAE in modeling open air systems. While Data-Driven Dynamics can compensate for some deficiencies in the physical equations, it still cannot match the effectiveness of direct physical modeling.
@@ 4.4 Case Study
Figure 7: Visualization of $\beta$, where negative and positive values separately indicate dissipation and generation. The arrow direction represents wind direction, and the length reflects wind speed at each station.
Next, we visualize the $\beta$ values in BA-DAE under two scenarios on the Beijing dataset (see Fig.7). As described in Section 3.2.2, negative values represent $\text{PM}_{\text{2.5}}$ dissipation (blue colorbar), while positive values indicate $\text{PM}_{\text{2.5}}$ generation (red colorbar). We observe two distinct scenarios: In the first scenario (left side of Fig.7), with wind present, dissipation significantly increases at boundary stations in the wind direction (denoted by circles), resulting in larger negative $\beta$ values. Additionally, some boundary stations exhibit positive $\beta$ values, which is reasonable because the wind at these stations carries pollutants from outside the region. In other words, this can be considered as pollutant generation. In the second scenario (right side of Fig.7), with no wind, boundary stations exhibit greater dissipation due to forest absorption and diffusion beyond the boundary, leading to negative $\beta$ values at the outskirts of Beijing. Notably, the city center of Beijing exhibits higher $\beta$ values at 9 a.m., indicating pollutant generation. This is expected, as 9 a.m. coincides with Beijing’s morning rush hour, when a large number of vehicles are present in the area. This example demonstrates that our model successfully captures the increase in pollutants caused by vehicle emissions. Overall, by integrating physical equations, Air-DualODE demonstrates excellent interpretability.
5 Conclusion and Future work
We design Air-DualODE for air quality prediction at both the city and national levels. To the best of our knowledge, this is the first dual dynamics model that accounts for open air systems. We employ a dual-dynamics approach, modeling both physical (known) and data-driven (unknown) dynamics using Neural ODEs, and subsequently align and fuse these dynamics in the latent space. This effectively integrates physical knowledge with data-driven methods. Compared to previous physics-informed approaches, our model can handle air quality prediction on both city-level and national-level datasets. In the future, we plan to further explore Physics-informed DualODE and extend its application to more general spatiotemporal prediction tasks.

### v2

# Air Quality Prediction with Physics-Guided Dual Neural ODEs in Open Systems

@@ Abstract
Air pollution significantly threatens human health and ecosystems, necessitating effective air quality prediction to inform public policy. Traditional approaches are generally categorized into physics-based and data-driven models. Physics-based models usually struggle with high computational demands and closed-system assumptions, while data-driven models may overlook essential physical dynamics, confusing the capturing of spatiotemporal correlations. Although some physics-guided approaches combine the strengths of both models, they often face a mismatch between explicit physical equations and implicit learned representations. To address these challenges, we propose Air-DualODE, a novel physics-guided approach that integrates dual branches of Neural ODEs for air quality prediction. The first branch applies open-system physical equations to capture spatiotemporal dependencies for learning physics dynamics, while the second branch identifies the dependencies not addressed by the first in a fully data-driven way. These dual representations are temporally aligned and fused to enhance prediction accuracy. Our experimental results demonstrate that Air-DualODE achieves state-of-the-art performance in predicting pollutant concentrations across various spatial scales, thereby offering a promising solution for real-world air quality challenges. The code is available at: https://github.com/decisionintelligence/Air-DualODE.
@@ 1 Introduction
Traditional research on air quality prediction largely relies on physics-based and data-driven approaches, each of which has their own inherent limitations. Physics-based models make predictions by solving Ordinary or Partial Differential Equations (ODEs/PDEs), which capture the transport of pollutants through diffusion and advection, as illustrated in Fig.1. Although these models achieve high accuracy, they require substantial computational resources, especially when applied on large spatial scales. Conversely, data-driven models excel at uncovering dependencies within data (Cheng et al., 2023; Xu et al., 2024a), yet their lack of integration with physical principles can lead to incomplete or even erroneous representations of spatiotemporal relationships. Therefore, it is crucial to propose a hybrid model that effectively harnesses the strengths of both physics-based and data-driven approaches. However, the development of such a model presents two primary challenges.
@@
In this paper, we propose Air-DualODE, a Physics-guided Dual Neural ODEs that integrate physics dynamics and data-driven dynamics for open system’s Air Pollution Prediction. For the first challenge, we propose a discrete Boundary-Aware Diffusion-Advection Equation (BA-DAE) to appropriately model open air systems. Due to the uncertainties associated with pollutant generation (source) and dissipation (sink) in open air systems, pollutant changes involve significant uncertainty, but both processes are linked to individual stations. To capture these dynamics, BA-DAE introduces a linear correction term, breaking away from the unrealistic closed system assumption and providing a more accurate description of physical phenomena in open air systems. Additionally, we incorporate Data-Driven Dynamics to capture spatiotemporal dependencies that BA-DAE ignores, making our approach better suited for real-world open air systems. For example, Data-Driven Dynamics can learn some patterns of pollutants generation, which cannot be described by BA-DAE.
@@ 2.3 Related Work
Air Quality Prediction. Air quality prediction is a key objective in smart city. Existing approaches can be broadly categorized into two types: physics-based models and deep learning models. The former rely on domain knowledge to construct Ordinary and Partial Differential Equations (ODEs/PDEs) that describe air pollutant transport (Li et al., 2023; Daly & Zannetti, 2007). However, these models face two major limitations. They rely on closed-system assumptions, and their use of numerical methods, such as finite difference or finite element methods (Özişik et al., 2017; Ŝolín, 2005), becomes computationally expensive at larger spatial scales. This makes rapid prediction difficult for urban planning (Lai et al., 2023) and policy-making (Liu et al., 2025). Deep learning (Campos et al., 2023; 2024; Qiu et al., 2024; Chen et al., 2024; Qiu et al., 2025) methods have recently gained prominence by using historical data and auxiliary covariates, such as weather, to extract high-dimensional spatiotemporal representations for pollutant prediction. These models typically employ RNNs or Transformers for temporal representations and GNNs or CNNs for spatial dependencies (Yu et al., 2017; Shang et al., 2021; Zhao et al., 2023; Wu et al., 2024). Furthermore, there are models specifically designed for air pollutant prediction tasks (Yu et al., 2023; Liang et al., 2023). However, these deep learning approaches may overlook essential physical dynamics, leading to confused spatiotemporal correlations that could violate physical principles.
Physics Guided Deep Learning. Recently, several studies have explored incorporating physical knowledge into deep learning models (Karpatne et al., 2024). The most direct approach involves using physical knowledge as constraints in loss functions (Raissi et al., 2017; Shi et al., 2021), penalizing deviations from physical laws to encourage physically consistent patterns and improve generalizability (Greydanus et al., 2019; Choi et al., 2023b). However, these methods rely on highly accurate physical knowledge; otherwise, they may introduce inductive bias, limiting the model’s learning capacity. Another approach is to design hybrid frameworks that integrate scientific knowledge into specific modeling components (Bao et al., 2021; Xu et al., 2024b), such as incorporating physical equations into Neural ODEs (Hwang et al., 2021; Ji et al., 2022; Choi et al., 2023a; Hettige et al., 2024). These methods often require deep latent representations to align with physical meanings, forcing models to switch between abstract representations and physical quantities Nascimento et al. (2021); Verma et al. (2024).
@@ 3.2.2 Discrete Boundary-Aware Diffusion-Advection Equation for open system
In an open system, changes in the total amount of pollutants can be categorized into dissipation and generation. Pollutant dissipation refers to sinks, including transport out of the boundary through diffusion and advection, as well as absorption by forests or lakes. Pollutant generation refers to sources such as industrial emissions. As shown in Fig.4, the dissipation and generation at Point E are highly relevant to its current concentrations. Therefore, we incorporate two terms into Eq.3: the dissipation term $\bm{\beta^{-}}\odot X$ ($\bm{\beta^{-}}\in[-1,0]^{N\times 1}$) and the generation term $\bm{\beta^{+}}\odot X$ ($\bm{\beta^{+}}\in[0,+\infty)^{N\times 1}$). These terms can be simplified into an open system correction term: $\bm{\beta}\odot X$ ($\bm{\beta}\in[-1,+\infty)^{N\times 1}$). Boundary stations, experiencing more outward diffusion and advection, tend to have greater negative $\beta_{i}$. So, this equation is called as Discrete Boundary-Aware Diffusion-Advection Equation (BA-DAE):
$\mathbf{F^{P}}(X_{t};\mathbf{\Theta})=\frac{dX}{dt}=\bm{\alpha}\odot(-k\cdot L_{\text{diff}}X)+(1-\bm{\alpha})\odot(L_{\text{adv}}X)+\bm{\beta}\odot X.$ (6)
Here, $\mathbf{\Theta}$ represents learnable parameters used to approximate the graph Laplacian operators. The gate value $\bm{\alpha}\in\mathbb{R}^{N\times 1}$ is estimated by a linear layer ($\mathbb{R}^{2}\to\mathbb{R}$) based on each station’s diffusion and advection over time, representing the weighted influence of these two effects as they vary across time and space. The diffusion coefficient $k$ and the correction term’s coefficient $\bm{\beta}$ are estimated via the Coefficient Estimator, a Recurrent Neural Network. By solving BA-DAE with ODE solver from Chen et al. (2018), pollutant concentrations from $T+1$ to $T+\tau$ are obtained, and a temporal encoder maps these concentrations to the latent space for dynamics fusion, as illustrated in the Physics Dynamics part of Fig 2.
@@ 3.3 Data-Driven Neural ODE
where $\mathbf{\Phi}$ represents the learnable parameters in ODE function $\mathbf{F^{D}}$. To capture spatial correlations across different time steps within the dynamic system, a masked self-attention mechanism is incorporated into $\mathbf{F^{D}}$. Specifically, the adjacency matrix of $G$ is employed as a mask in the Spatial-MSA since intuitively, if no potential transport pathway exists between two stations, their representations should not be correlated. This approach not only improves computational efficiency but also enables the model to focus on relevant information, enhancing its effectiveness across various spatial granularities in prediction scenarios.
@@ 3.4 Dynamics Fusion
Decaying Temporal Constrastive Learning. Before combining the physics and data-driven dynamics, it is essential to align their latent representations, $Z^{P}_{T+1:T+\tau}$ and $Z^{D}_{T+1:T+\tau}$, to resolve inconsistencies and enable effective fusion. Since both dynamics operate synchronously, three key intuitions are considered: 1) latent representations at the same time step are more similar than those at different time steps, 2) adjacent time steps are more similar than distant ones, and 3) representations from the same dynamics are more similar than those from different dynamics. To satisfy the three intuitions, we propose Decaying Temporal Contrastive Learning (Decay-TCL). It computes a decaying weight $w(t,s)$ based on timestamp differences and dynamics type to regulate the similarity between latent representations. The hyperparameters $\lambda_{1}$ and $\lambda_{2}$ are used to differentiate the distinct dynamics, and they must satisfy $\lambda_{1}>\lambda_{2}$ to fulfill third intuition.
@@
GNN Fusion. After temporal alignment, the two latent representations are concatenated as $Z^{(0)}_{t}=\text{concat}(Z^{P}_{t},Z^{D}_{t})$ at each timestamp. Since the influence of stations is distance-dependent, using MLP structures may introduce noises from distant sites, complicating the learning process thus reducing accuracy (see Appendix A.16). To address this, GNNs are employed based on the geospatial graph $G$ to Dynamics Fusion: $Z^{(n)}_{t}=\text{GNNs}(Z^{(0)}_{t})$, where $n$ represents the GNN layers. The Decoder then transforms the latent representations $Z^{(n)}_{T+1:T+\tau}$ into predicted pollutant concentrations $\hat{X}_{T+1:T+\tau}$. We jointly optimize the prediction loss and the loss functions in Eq.12, with a hyperparameter $\gamma$ to balance them:
$\mathcal{L}=\frac{1}{\tau}\sum_{t=T+1}^{T+\tau}\|X_{t}-\hat{X}_{t}\|+\gamma\mathcal{L}_{\text{tcl}}.\vspace{-5mm}$ (13)
@@ 4.1 Experiment Setting
Baselines. We evaluate Air-DualODE against baselines of four categories: 1) Classical Methods: Historical Average (HA) and Vector Auto-Regression (VAR). 2) Differential Equation Network-Based Methods: Latent-ODE  (Chen et al., 2018), ODE-RNN (Rubanova et al., 2019) and ODE-LSTM  (Lechner & Hasani, 2020). 3) Spatio-temporal Deep Learning Methods: DCRNN (Li et al., 2017), STGCN (Yu et al., 2017), ASTGCN  (Guo et al., 2019), GTS (Shang et al., 2021), MTSF-DG  (Zhao et al., 2023), PM2.5-GNN (Wang et al., 2020), and AirFormer (Liang et al., 2023). 4) Physics-guided Neural Networks: AirPhyNet (Hettige et al., 2024).
@@ 4.2 Performance Comparison
From Table 1 we can also observe the following: 1) deep learning approaches outperform classical methods such as HA and VAR, demonstrating their superior ability to capture complex spatiotemporal correlations. 2) Spatiotemporal deep learning models, originally developed for traffic flow forecasting, such as ASTGCN and GTS, exhibit strong adaptability to air quality forecasting, producing results comparable to SOTA air quality prediction methods like AirFormer and PM25GNN. 3) Approaches based on Neural DE networks also exhibit competitive performance compared to spatiotemporal deep learning methods, as they capture the ODE function in latent space, which is crucial for accurate air quality prediction. Overall, these results suggest the effectiveness of Air-DualODE in achieving precise air quality predictions while leveraging the physics-based domain knowledge of air pollutant transport. More Discussions about sudden change evaluation can be found in Appendix A.17.
@@ 4.3 Ablation Study
Effect of Physical knowledge. To further demonstrate the impact of incorporating physical knowledge, we compare Air-DualODE with three other closed-system variations: a) Diff-c, which only considers the diffusion process, b) Adv-c, which only considers the advection process, and c) DAE-c, which accounts for both processes. Additionally, we compare these with d) BA-DAE-o, namely Air-DualODE, as proposed in Section 3.2.2 for open air systems. Fig.5 shows that integrated physical knowledge is crucial, and inadequate modeling leads to reduced performance. Consistent with AirPhyNet, DAE-c outperforms Diff-c and Adv-c, but its inability to model dissipation and generation makes it less effective than BA-DAE-o. This underscores the effectiveness of BA-DAE in modeling open air systems. While Data-Driven Dynamics can compensate for some deficiencies in the physical equations, it still cannot match the effectiveness of direct physical modeling.
@@ 4.4 Case Study
Figure 7: Visualization of $\bm{\beta}$, where negative and positive values separately indicate dissipation and generation. The arrow direction represents wind direction, and the length reflects wind speed at each station.
Next, we visualize the $\bm{\beta}$ in BA-DAE under two scenarios on the Beijing dataset (see Fig.7). As described in Section 3.2.2, negative values represent $\text{PM}_{\text{2.5}}$ dissipation (blue colorbar), while positive values indicate $\text{PM}_{\text{2.5}}$ generation (red colorbar). We observe two distinct scenarios: In the first scenario (left side of Fig.7), with wind present, dissipation significantly increases at boundary stations in the wind direction (denoted by circles), resulting in larger negative $\beta_{i}$. Additionally, some boundary stations exhibit positive $\beta_{i}$, which is reasonable because the wind at these stations carries pollutants from outside the region. In other words, this can be considered as pollutant generation. In the second scenario (right side of Fig.7), with no wind, boundary stations exhibit greater dissipation due to forest absorption and diffusion beyond the boundary, leading to negative $\beta_{i}$ at the outskirts of Beijing. Notably, the city center of Beijing exhibits higher $\beta_{i}$ at 9 a.m., indicating pollutant generation. This is expected, as 9 a.m. coincides with Beijing’s morning rush hour, when a large number of vehicles are present in the area. This example demonstrates that our model successfully captures the increase in pollutants caused by vehicle emissions. Overall, by integrating physical equations, Air-DualODE demonstrates excellent interpretability.
5 Conclusion and future works
We design Air-DualODE to predict air quality at both city and national levels. Our method employs a dual-dynamics approach, leveraging Neural ODEs to model physical (known) and data-driven (unknown) dynamics, which are subsequently aligned and fused in the latent space. By integrating physical equations into deep learning, Air-DualODE allows existing physical knowledge to enhance the model’s learning, inference, and ultimately, its prediction accuracy. While the results are promising, predicting sudden changes remains a significant challenge in air quality forecasting. Therefore, Designing specific mechanisms to better handle sudden changes is an important and promising research direction. Moreover, we plan to further explore Physics-guided DualODE and extend its application to more general spatiotemporal prediction tasks.
Acknowledgement
This work is partially supported by National Natural Science Foundation of China No.62372179. Yuxuan Liang‘s work is supported by the National Natural Science Foundation of China No.62402414.